Find 3 consecutive integers such that two times the sum of all three is 3 times the product of the larger two
Answer by jorel555(1290)

Let n be your first integer. Then your other integers are n+1 and n+2. So:
2(n+n+1+n+2)=3(n+1)(n+2)
6n+6=3nē+9n+6
3nē+3n=0
3nē=-3n
n=-1
n+1=0
n+2=1. ☺☺☺☺